Hi! I was wondering if you could help me with the following problem:
Find the slope of the line perpendicular to -3x+2y=15
If you could help that'd be great!
Thanks.

Answer by dolly(163)

Given line is -3x + 2y = 15
==> 2y = 3x + 15
==> y = 3/2 x + 15/2
This is of the form y = mx + c where m is the slope.
So slope of this line is 3/2
Therefore the slope of the perpendicular line is -1/(3/2){as product of the slopes of perpendicular ines is -1 always]
= -2/3
Thus the slope of the line perpendicular to the given line is -2/3